OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3)

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

find x ?

OpenStudy (anonymous):

distribute and simplify both sides first....can you tell me what you get after that?

OpenStudy (anonymous):

\[-2 (-3+x)\geq 2-x \]Expand the LHS.\[6-2 x\geq 2-x \]Subtract 6 from each side.\[-6+6-2 x\geq 2-x-6 \]\[-2 x\geq -4-x \]Add 2x to each side.\[2 x-2 x\geq 2 x-4-x \]\[0\geq -4+x \]Add 4 to each side.\[4+0\geq -4+x+4 \]\[4\geq x \]\[x\leq 4 \]

OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) you are going to multiply -2 with x and 3 then multiply 5 with x and 3 then you should get -2x - 6 and 5x+15 then combine like terms

OpenStudy (anonymous):

@chikitagee 5-(x+3) is not equal to 5*(x+3)

OpenStudy (anonymous):

then what is it @robtobey

OpenStudy (anonymous):

≥ 5 – (x + 3) -> ≥ 5 – x - 3)